A galvanometer of resistance 50 ohm which gives a full scale deflection for 1mA is to be adapted to measure currents of up to 5A if this resistor is made of a material of cross sectional area 4×10^-4 and resistivity 2×10^-6. Calculate it's length
1
Expert's answer
2020-04-14T09:42:12-0400

"I(R)=I\u2212I(r)=5A\u22120.001A=4.999A;I(R)\\times R=r\\times I(r);R=\\frac{r\\times I(r)}{I(R)}=\\frac{50\\times 0.001}{4.999}=0.01Ohm;L=\\frac{4\\times 10^{-4} \\times 0.01}{2\\times 10^{-6}} = 2m I(R)\u00d7R=r\u00d7I(r);R= \nI(R)\nr\u00d7I(r)\n\u200b\t\n = \n4.999\n50\u00d70.001\n\u200b\t\n =0.01Ohm;L= \n2\u00d710 \n\u22126\n \n4\u00d710 \n\u22124\n \u00d70.01\n\u200b\t\n =2m"